OpenStudy (anonymous):

!!!!!PLEASE SOMEONE HELP ME IM DESPERATE WILL GIVE MEDAL!!!!! Solve the following equation by completing the square x^2 - 10x = 13

OpenStudy (anonymous):

subtract 13 it appears to be a quadratic, so use the quad formula to solve.

OpenStudy (anonymous):

okay so i get x^2 - 10x -13 =0?

OpenStudy (anonymous):

yes so use this: \[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

a=1 b=10 c=-13

OpenStudy (anonymous):

No thats not what the equation is asking. Its asking to complete the square

OpenStudy (anonymous):

OH!! yes, one moment please

OpenStudy (anonymous):

do you know how to do it then?

OpenStudy (anonymous):

@Samidancer16 ?

OpenStudy (anonymous):

oky

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Now we can solve a Quadratic Equation in 5 steps: Step 1 Divide all terms by a (the coefficient of x2). Step 2 Move the number term (c/a) to the right side of the equation. Step 3 Complete the square on the left side of the equation and balance this by adding the same value to the right side of the equation. We now have something that looks like (x + p)2 = q, which can be solved rather easily: Step 4 Take the square root on both sides of the equation. Step 5 Subtract the number that remains on the left side of the equation to find x.

OpenStudy (anonymous):

ok x^2 - 10x -13 =0 x^2 - 10x -13+ =25?

OpenStudy (anonymous):

x^2 - 10x -13+25 =25?

OpenStudy (anonymous):

is that the starting equation?

OpenStudy (anonymous):

ye now wut?

OpenStudy (anonymous):

wut do i do now?

OpenStudy (anonymous):

\[x=5\pm \sqrt{38}\]